# LLM-Based Detection of Tangled Code Changes for Higher-Quality Method-Level Bug Datasets

### v2

@@ 3.2. Prompt Preparation
The design of our prompt structure draws on recent advancements in prompt engineering that demonstrate the effectiveness of decomposing complex instructions into clearly scoped components (Brown et al., 2020; Wei et al., 2023; Zhou et al., 2023). As prior work has shown, LLMs are highly sensitive to prompt formulation, and performance can vary substantially based on how the task, instructions, and examples are framed (Liu et al., 2023). To maximize reliability and adaptability, we designed a modular prompt template consisting of three core components, aligned with prior studies (Amatriain, 2024; Giray, 2023), as illustrated in Figure 1.
Instructions. This component defines the LLM’s role and expectations. It is composed of four parts: (i) a role-based persona (e.g., “You are a Git commit review assistant”), which helps constrain the model’s responses by grounding its behavior and improves performance (Hu and Collier, 2024); (ii) a high-level task description (Amatriain, 2024); (iii) step-by-step behavioral guidance to structure the model’s reasoning (Wei et al., 2023); and (iv) explicit output formatting requirements to reduce ambiguity and improve parsability (Liu et al., 2024a).
Few-Shot Examples. Few-shot prompting has been shown to significantly improve performance by enabling in-context learning, especially in code and software engineering tasks (Chen et al., 2021). For few-shot examples, as shown in Figure 2, we selected a representative bug-fix commit containing one diff genuinely related to a bug fix and another unrelated change, a real instance of tangled code changes. These two fixed examples were consistently utilized to prime the LLMs in all experiments employing few-shot learning. These examples were selected to represent contrasting scenarios, one clearly indicative of a bug-fix and the other representing a refactor or performance-related change that does not qualify as a bug-fix, although both of the examples come from the same commit.
@@
Figure 2. Commit message and two methods modified in the same commit: the message indicates a bug fix, but only the method on the right contains the bug-fix change, while the method on the left is a non-bug-related tangled change. Used as few-shot examples in our prompts.
[…]
diff
LLMs like GPT-4o and Gemini-2.0-Flash often exhibit significant performance degradation as prompt length increases, due to limitations in handling extended context (Levy et al., 2024; Zhang et al., 2024). Additionally, longer prompts incur higher computational and monetary costs, making token efficiency a practical concern. Since method-level diffs can vary greatly in length and the input query, over which we have no control, is the primary driver of prompt size, we deliberately restricted the prompt to a single example per class (Buggy and NotBuggy). This prompting approach helps maintain prompt lengths within manageable token limits across different model families. By limiting the number of in-context examples, we strike a balance between performance and resource usage.
@@
As shown in Figure 2, for the negative example (NotBuggy), we selected the code diff that merely replaced constant values with equivalent boolean literals, without altering the program’s logic or behavior. In contrast, the other code diff (Buggy) addressed an issue in the byte-to-hexadecimal conversion logic by correctly handling signed bytes using bit masking, rectifying a functional error.
This strategy aligns with established practices in few-shot prompting  (Brown et al., 2020; Min et al., 2022), where concise and diverse instances are used to anchor the model’s reasoning prior to inference. In our setup, the examples provide contextual grounding by showcasing typical linguistic and structural patterns present in both buggy and non-buggy commits.
@@
It is important to note that each research question required slight variations in prompt configuration, e.g., enabling or disabling few-shot examples, toggling chain-of-thought reasoning, or isolating input modalities (diff-only vs. diff + message).
@@ 3.3. Selecting and Querying LLMs
Selecting LLMs. We selected three state-of-the-art proprietary LLMs for evaluation: GPT-4o-mini, GPT-4o, and Gemini-2.0-flash. These models were chosen based on a combination of practical, technical, and empirical factors, including competitive performance on recent code-related benchmarks (Google DeepMind, 2024; OpenAI, 2025; Batista et al., 2024; Bruni et al., 2025; Demirtaş et al., 2025), robust API availability, and cost-effectiveness for large-scale experimentation.
GPT-4o-mini: A cost-efficient, lightweight model from OpenAI with reduced parameter (OpenAI, 2025). This LLM is used in various software engineering research, such as code generation and analysis (Bruni et al., 2025; Demirtaş et al., 2025).
GPT-4o: OpenAI’s full-scale model, designed for maximum performance. This model is proven to be more robust than its smaller variant gpt-4o-mini (OpenAI, 2025) in code-related tasks (Bruni et al., 2025; Demirtaş et al., 2025).
Gemini-2.0-Flash: A high-performing model from Google’s Gemini series (Google DeepMind, 2024), delivering competitive performance to the GPT series (Islam et al., 2025) and used in various code-related tasks (Batista et al., 2024).
Due to computational limitations, we did not include open-source models. However, according to many previous studies, the selected proprietary models generally perform better than the open-source counterparts on multiple standard benchmarks, including code-related tasks (Haroon et al., 2025; Islam et al., 2025; Bae et al., 2024). If our selected LLMs cannot solve the tangling problem, it is unlikely that other models can. Their performance is likely to be the upper bound for all LLMs. However, we acknowledge that more experiments with other models can produce interesting results.
@@ 4.1. RQ1: Performance of zero-shot LLMs using code diff and commit message
Diff-Only Detection. The LLM is provided only with the code diff and instructed to assess whether the change represents a bug fix (Buggy) or not (NotBuggy). The prompt is adapted from Figure 1 by removing all references to commit messages. A representative fragment of the prompt is: ”… Given a Java source code diff, analyze it to determine if the changes are related to a bug fix. Examine the modifications carefully, looking for error-handling …”. The output formatting is also modified to: ”Output only a single word: ’Buggy’ if the changes indicate a bug fix, otherwise ’NotBuggy’.”, since the format in Figure 1 is intended for chain-of-thought.
Diff+Message Detection. The prompt includes both the code diff and its commit message, instructing the model to assess their semantic alignment to identify bug-fixing intent. This prompt follows Figure 1, except for adopting the same output format constraint as in the Diff-Only Detection setting.
Table 1 compares LLM performance under two input scenarios: Diff-only and Diff+Message. In the Diff-only setting, gpt-4o-mini achieves the highest performance across all metrics, with an F1 score of 0.747. When commit messages are included, all models show substantial improvements. gpt-4o yields the best overall results, attaining the highest accuracy (0.874), F1 score (0.877), and MCC (0.747). Although gemini-2.0-flash achieves the highest recall (0.988), its low precision (0.717) leads to a reduced F1 score (0.831). These findings indicate that incorporating commit messages alongside diffs consistently enhances defect detection performance.
@@
Diff-only gpt-4o-mini 0.704 0.660 0.860 0.747 0.426
@@
Diff + Message gpt-4o-mini 0.822 0.764 0.941 0.843 0.661
@@
Summary of RQ1: LLMs can effectively detect tangled code changes using both code diffs and commit messages, with combined input yielding better results than code diff alone.
@@ 4.2. RQ2: Performance of different prompting techniques
The effectiveness of LLMs highly depends on the input prompt, and the success of different prompting techniques is proven in various studies (Wei et al., 2022; Brown et al., 2020; Wei et al., 2023). This research question investigates how different prompting techniques, namely few-shot learning, chain-of-thought prompting, and a hybrid approach combining both, impact the classification performance of LLMs.
Based on the findings from RQ1, where the combination of commit message and code diff outperformed the diff-only setting, we use this input setting for RQ2. Additionally, since gpt-4o significantly outperformed gpt-4o-mini, we exclude the latter from further experiments. This decision also supports cost efficiency, as the prompting techniques used in RQ2 involve substantial token consumption in both input and output.
We implemented the following prompting configurations:
@@
Table 2 presents the performance of different prompting strategies across LLMs. While gemini-2.0-flash with few-shot achieves the highest recall (0.989), it suffers from low precision (0.723). Conversely, gpt-4o with few-shot yields the highest precision (0.903) but lower recall (0.833). The combined few-shot and chain-of-thought prompting delivers the most balanced performance, with gpt-4o attaining the highest accuracy (0.880), F1 score (0.883), and MCC (0.760). These findings highlight the effectiveness of combining reasoning and exemplars for tangled change detection.
Table 2. Performance comparison of prompting techniques for tangled change detection (gemini refers to gemini-2.0-flash). Metrics are calculated by targeting the buggy samples (pos_label = ”Buggy”).
@@
Few-shot gpt-4o 0.870 0.903 0.833 0.866 0.742
gemini 0.802 0.723 0.989 0.835 0.649
Chain-of-thought gpt-4o 0.827 0.770 0.940 0.846 0.669
gemini 0.765 0.688 0.984 0.810 0.587
Chain-of-thought + Few-shot gpt-4o 0.880 0.871 0.896 0.883 0.760
gemini 0.807 0.731 0.979 0.837 0.651
A comparison between Table 1 and 2 reveals that few-shot prompting alone does not improve performance over zero-shot prompting in RQ1. In fact, chain-of-thought prompting on its own performs worse. However, combining few-shot prompting with chain-of-thought yields the best performance, surpassing all other configurations. This pattern aligns with findings from prior research, which suggests that few-shot prompting may cause models to overfit to the format of examples rather than the task itself, while zero-shot chain-of-thought prompting can amplify bias and lead to suboptimal results (Kojima et al., [n. d.]; Shaikh et al., 2023). Our findings are consistent with earlier work showing that the combination of few-shot and chain-of-thought prompting generally outperforms zero-shot methods (Kojima et al., [n. d.]).
Table 3. Predictions on a semantically misleading commit.
@@
To better understand why the chain of thought technique improves model performance in the few shot setting, we analyze a representative example involving semantic ambiguity between the commit message and the code modifications from the PMD project. [^2] https://github.com/pmd/pmd As shown in Table 3, the commit message clearly describes a bug fix involving two files, PMD.java and ImmutableFieldRule.java. In this commit [^3] https://github.com/pmd/pmd/commit/a405d23dfb9e574e2b2ef23f1f45d548a738ed3b, multiple methods are modified, including initializedInConstructor() from ImmutableFieldRule.java, which only includes variable renaming (e.g., occurance to occ) without affecting the program’s logic or behavior. No structural fix is present that addresses the reported issue of static field misclassification. Despite this, both the zero-shot and few-shot models incorrectly classify the change as Buggy, influenced by the commit message. In contrast, the chain-of-thought technique, with or without few-shot prompting, correctly identifies the change as NotBuggy, consistent with the manually assigned label.
The reasoning produced by the chain-of-thought prompt illustrates a more grounded assessment: The commit fixes a bug associated with the class ImmutableFieldRule to prevent false positives for static fields. However, in the provided diff, the only change is renaming the variable occurance to occ, which is a simple refactor for readability or consistency purposes. This modification itself doesn’t address the logic or behavior of the code in terms of resolving the stated bug. The commit message also mentions a version number fix in PMD.java, but no relevant changes are shown in this diff. The diff provided does not include any alterations that directly fix the bug mentioned regarding false positives or affect the logic related to the bug description. Since the diff doesn’t contain changes indicating any error-handling updates, logical corrections, exception-handling improvements, or other signs of bug-related changes associated with addressing false positives for static fields, the review of the diff itself does not point towards a bug fix. Thus, based on the diff provided, the conclusion for this specific change is **NotBuggy**.
This case highlights the effectiveness of chain-of-thought prompting in systematically analyzing complex commit messages that reference multiple issues. While such messages offer valuable context, they can also introduce ambiguity. The step-by-step reasoning decomposed the message, independently assessed each part against the code diff, and correctly identified the lack of structural or semantic alignment, ultimately concluding that the change was not a bug fix.
Summary of RQ2: Combining chain-of-thought reasoning with few-shot examples significantly improves the effectiveness of LLMs for detecting tangled code changes, outperforming each prompting strategy. Chain-of-thought reasoning excels by encouraging the model to critically assess the semantic alignment between commit messages and code diffs, as shown in the case study, where it correctly identified non-buggy changes despite a misleading bug-fix commit message.

### v3

@@ 3.2. Prompt Preparation
The design of our prompt structure draws on recent advancements in prompt engineering that demonstrate the effectiveness of decomposing complex instructions into clearly scoped components (Brown et al., 2020; wei_chain--thought_2023; zhou_least--most_2023). As prior work has shown, LLMs are highly sensitive to prompt formulation, and performance can vary substantially based on how the task, instructions, and examples are framed (Liu et al., 2023). To maximize reliability and adaptability, we designed a modular prompt template consisting of three core components, aligned with prior studies (Amatriain, 2024; Giray, 2023), as illustrated in Figure 1.
@@
Figure 2. Commit message and two methods modified in the same commit: the message indicates a bug fix, but only the method on the right contains the bug-fix change, while the method on the left is a non-bug-related tangled change. This was used as few-shot examples in our prompts.:
Instructions. This component defines the LLM’s role and behavior. It is composed of four parts: (i) a role-based persona (e.g., “You are a Git commit review assistant”), which helps constrain the model’s responses by grounding its behavior and improves performance (Hu and Collier, 2024); (ii) a high-level task description (Amatriain, 2024); (iii) step-by-step behavioral guidance to structure the model’s reasoning (wei_chain--thought_2023); and (iv) explicit output formatting requirements to reduce ambiguity and improve parsability (Liu et al., 2024).
@@
Few-shot Examples. Few-shot prompting has been shown to significantly improve performance by enabling in-context learning, especially in code and software engineering tasks (zheng2024few). For few-shot examples, as shown in Figure 2, we selected a representative bug-fix commit containing one diff genuinely related to a bug fix and another unrelated change, a real instance of tangled code changes. For the negative example (NotBuggy), we selected the code diff that merely replaced constant values with equivalent boolean literals, without altering the program’s logic or behavior. In contrast, the other code diff (Buggy) addressed an issue in the byte-to-hexadecimal conversion logic by correctly handling signed bytes using bit masking, rectifying a functional error. These two fixed examples were consistently utilized to prime the LLMs in all experiments employing few-shot learning. These examples were selected to represent contrasting scenarios, one clearly indicative of a bug-fix and the other representing a refactor or performance-related change that does not qualify as a bug-fix, although both of the examples come from the same commit.
LLMs often exhibit significant performance degradation as prompt length increases, due to limitations in handling extended context (Levy et al., 2024; zhang_inftybench_2024). Additionally, longer prompts incur higher computational and monetary costs, making token efficiency a practical concern. Since method-level diffs can vary greatly in length and the input query, over which we have no control, is the primary driver of prompt size, we deliberately restricted the prompt to a single example per class (Buggy and NotBuggy). This prompting approach helps maintain prompt lengths within manageable token limits across different model families. By limiting the number of in-context examples, we strike a balance between performance and resource usage. This strategy aligns with established practices in few-shot prompting  (Brown et al., 2020), where concise and diverse instances are used to anchor the model’s reasoning prior to inference.
@@
Each research question required slight variations in prompt configuration, e.g., enabling or disabling few-shot examples, toggling chain-of-thought reasoning, or isolating input modalities (diff-only vs. diff + message). All corresponding prompt templates are included in our replication package, located within the prompts directory.
@@ 3.3. Selecting and Querying LLMs
Selecting LLMs. For RQ1 and RQ2, we selected six LLMs: four proprietary, GPT-4o-mini, GPT-4o, GPT-5, and Gemini-2.0-Flash and two open-source, GPT-OSS-20B and GPT-OSS-120B. The proprietary LLMs were chosen based on their performance on recent code-related benchmarks (Google DeepMind, 2024; OpenAI, 2025; Batista et al., 2024; Bruni et al., 2025), robust API support, and cost-effectiveness for large-scale experimentation. GPT-4o serves as OpenAI’s full-scale model offering superior robustness over its smaller variant, GPT-4o-mini (OpenAI, 2025; Bruni et al., 2025), while GPT-5 represents the latest generation of OpenAI’s LLMs. Gemini-2.0-Flash, from Google (Google DeepMind, 2024), delivers competitive performance with the GPT series and has demonstrated effectiveness in code-related tasks (Batista et al., 2024).
For open-source alternatives and to ensure reproducibility and accessibility, we choose GPT-OSS-20B (requiring  16 GB GPU memory) and GPT-OSS-120B (runnable on a single 80 GB GPU) (Agarwal et al., 2025). Despite their moderate scale, these models deliver reasoning and code-understanding performance comparable to proprietary baselines (Agarwal et al., 2025; Bi et al., 2025), balancing efficiency and capability.
For RQ3, we employed four models for embedding generation: one proprietary and three open-source. The proprietary model, Gemini’s text-embedding-004 offers high-quality embeddings optimized for classification tasks (Lee et al., 2024). Open-source models, CodeBERT (Feng et al., 2020) and GraphCodeBERT (Guo et al., 2020) are widely adopted for code representation learning (Ma et al., 2024), while EmbeddingGemma-300m, a recent lightweight model from Google DeepMind, is designed for low-latency, high-throughput applications (vera2025embeddinggemma). The models used in RQ1 and RQ2 were not included in RQ3 because they are decoder-only generative models optimized for text generation, whereas RQ3 required encoder-based architectures specialized for producing fixed-length embeddings suitable for classification tasks.
In summary, proprietary models were selected for their state-of-the-art performance across benchmarks, including code-related tasks (Batista et al., 2024; Bae et al., 2024), while open-source models were selected for their competitive accuracy and efficiency relative to larger open-source alternatives.
@@ 3.4. Evaluation Metrics
3.5. Experimental Setup
All experiments involving open-source LLMs were conducted on a High-Performance Computing (HPC) cluster equipped with Intel 8570 CPUs (2.1 GHz) and multiple NVIDIA H100 GPUs. The number of CPUs, GPUs and the memory allocation varied by model, depending on resource availability and scheduling within the cluster.
@@ 4.1. RQ1: Performance of zero-shot LLMs using code diff and commit message
Diff-Only. The LLM is provided only with the code diff and instructed to assess whether the change represents a bug fix (Buggy) or not (NotBuggy). The prompt is adapted from Figure 1, with all references to commit messages and few-shot examples removed, and the instruction modified to require a single-word response.
Diff+Message. The prompt includes both the code diff and its commit message, instructing the model to assess their semantic alignment to identify bug-fixing intent. This prompt follows the Diff-Only Detection setting, except that all references to commit messages are retained as in Figure 1.
Table 1 compares the performance of the evaluated LLMs under two input settings: Diff-only and Diff+Message. In the Diff-only scenario, gpt-5 achieves the best overall performance, yielding the highest F1-score (0.767) and MCC (0.490). When commit messages are included, all models show substantial improvements. The gpt-5 again attains the highest F1-score (0.879), while gpt-4o provides the best accuracy (0.874), precision (0.868) and MCC (0.747). Although gpt-5 achieves the top F1-score, its high precision–recall gap indicates less balanced performance, whereas gpt-4o offers the most consistent results across all metrics.
The gemini-2.0-flash and the gpt-oss models achieve high recall (>0.9>0.9) but low precision (<0.76<0.76) resulting in reduced F1-scores. The gpt-oss models perform remarkably well in Diff+Message scenario despite their smaller size narrowing the gap with proprietary models (within  3–4% of gpt-5’s F1-score), likely benefiting from their unique dual-channel chat format (Agarwal et al., 2025), where an internal “analysis” channel generates intermediate reasoning (chain-of-thought) before producing the answer on ”final” channel. Overall, these results demonstrate that integrating commit messages with diffs significantly improves performance, and that open-source LLMs are also effective despite their smaller sizes.
@@
Diff-only gpt-5 0.742 0.709 0.834 0.767 0.490
gpt-4o-mini 0.704 0.660 0.860 0.747 0.426
@@
gpt-oss-20b 0.663 0.615 0.900 0.730 0.364
gpt-oss-120b 0.667 0.619 0.900 0.733 0.372
Diff
Message
gpt-5 0.871 0.839 0.924 0.879 0.746
gpt-4o-mini 0.822 0.764 0.941 0.843 0.661
@@
gpt-oss-20b 0.820 0.758 0.950 0.843 0.662
gpt-oss-120b 0.827 0.766 0.948 0.847 0.671
Summary of RQ1: Both proprietary and open-source LLMs can effectively detect tangled code changes using code diffs, with or without commit messages, though combining both inputs yields better results than code diffs alone.
@@ 4.2. RQ2: Performance of different prompting techniques
The effectiveness of LLMs highly depends on the input prompt, and the success of different prompting techniques is proven in various studies (wei_finetuned_2022; Brown et al., 2020; wei_chain--thought_2023). This research question investigates how different prompting techniques, namely few-shot learning, chain-of-thought prompting, and a hybrid approach combining both, impact the classification performance of LLMs.
@@
Based on the findings from RQ1, we use the combination of commit message and code diff for RQ2. Additionally, we exclude gpt-4o-mini from further experiments, as gpt-4o significantly outperformed this small model. This decision also supports cost efficiency, as the prompting techniques used in RQ2 involve substantial token consumption in both input and output.
Table 2 presents the performance of different prompting strategies across LLMs. While gpt-4o with few-shot achieves the highest precision (0.903) but lower recall (0.833), other models exhibit the opposite trend. In this setting, gpt-5 attains the best overall performance with an F1-score of 0.884. For chain-of-thought prompting, performance generally decreases across models, though gpt-5 again achieves the highest F1-score (0.872). When combining few-shot and chain-of-thought prompting, gpt-4o delivers the best results across all metrics except recall, which is highest for gemini-2.0-flash.
Table 2. Performance comparison of different prompting techniques. Metrics are calculated by targeting the buggy samples (pos_label = ”Buggy”).
@@
Few-shot gpt-5 0.876 0.842 0.931 0.884 0.757
gpt-4o 0.870 0.903 0.833 0.866 0.742
gemini-2.0-flash 0.802 0.723 0.989 0.835 0.649
gpt-oss-20b 0.837 0.775 0.958 0.857 0.694
gpt-oss-120b 0.819 0.753 0.959 0.843 0.663
Chain
of
thought
gpt-5 0.863 0.829 0.920 0.872 0.730
gpt-4o 0.827 0.770 0.940 0.846 0.669
gemini-2.0-flash 0.765 0.688 0.984 0.810 0.587
gpt-oss-20b 0.829 0.770 0.946 0.849 0.676
gpt-oss-120b 0.829 0.765 0.956 0.850 0.679
Chain
of
thought
Few-shot
gpt-5 0.863 0.829 0.921 0.873 0.731
gpt-4o 0.880 0.871 0.896 0.883 0.760
gemini-2.0-flash 0.807 0.731 0.979 0.837 0.651
gpt-oss-20b 0.844 0.795 0.932 0.858 0.697
gpt-oss-120b 0.816 0.752 0.952 0.840 0.655
Among all prompting strategies, gpt-5 with few-shot yields the top F1-score, while gpt-4o attains the highest accuracy, precision, and MCC with chain-of-thought+few-shot. Considering balanced performance across metrics, particularly the small gap between precision and recall (0.025), gpt-4o with chain-of-thought+few-shot offers the most balanced results, with an F1-score (0.883) nearly identical to the best (0.884). Open-source models also benefit consistently, though modestly, from advanced prompting; notably, gpt-oss-20b remains competitive, often matching or surpassing gpt-oss-120b.
A comparison between Table 1 and Table 2 reveals that few-shot prompting improves performance primarily for gpt-5 and slightly for gpt-oss-20b, while other models show marginal or decreased performance. Notably, chain-of-thought alone results in poorer performance across most models. However, combining few-shot prompting with chain-of-thought yields the most balanced and consistent outcomes overall, as observed with gpt-4o. This trend aligns with prior research suggesting that few-shot prompting can lead models to overfit to example formats rather than the task itself, while zero-shot chain-of-thought prompting may amplify biases and degrade reasoning quality (Kojima et al., 2022; shaikh_second_2023). Our findings thus corroborate earlier studies indicating that integrating few-shot and chain-of-thought generally outperforms zero-shot approaches (Kojima et al., 2022).
Table 3. Predictions on a semantically misleading commit using chain-of-thought + few-shot prompting with gpt-4o.
@@
To better understand why the chain of thought technique improves model performance in the few shot setting with gpt-4o, we analyze a representative example involving semantic ambiguity between the commit message and the code modifications from the PMD  [^3] https://github.com/pmd/pmd project. As shown in Table 3, the commit message clearly describes a bug fix involving two files, PMD.java and ImmutableFieldRule.java. In this commit [^4] https://github.com/pmd/pmd/commit/a405d23dfb9e574e2b2ef23f1f45d548a738ed3b, multiple methods are modified, including initializedInConstructor() from ImmutableFieldRule.java, which only includes variable renaming (e.g., occurance to occ) without affecting the program’s logic or behavior. No structural fix is present that addresses the reported issue. Despite this, both the zero-shot and few-shot models incorrectly classify the change as Buggy, influenced by the commit message. In contrast, the chain-of-thought technique, with or without few-shot prompting, correctly identifies the change as NotBuggy, consistent with the manually assigned label.
The reasoning produced by the chain-of-thought prompt illustrates a more grounded assessment. An important part of the reasoning states: ”. . . However, in the provided diff, the only change is renaming the variable occurance to occ, which is a simple refactor for readability or consistency purposes. This modification itself doesn’t address the logic or behavior of the code in terms of resolving the stated bug. The commit message also mentions a version number fix in PMD.java, but no relevant changes are shown in this diff. . . .”
This case highlights the effectiveness of chain-of-thought in systematically analyzing complex commit messages that reference multiple issues. While such messages offer valuable context, they can also introduce ambiguity. The step-by-step reasoning decomposed the message, independently assessed each part against the code diff, and correctly identified the lack of structural or semantic alignment, ultimately concluding that the change was not a bug fix.
Summary of RQ2: Combining chain-of-thought reasoning with few-shot prompting yields the most balanced and consistent performance across models. While few-shot prompting alone benefits mainly gpt-5, and chain-of-thought alone often reduces accuracy, their combination enhances reasoning stability and mitigates overfitting. This hybrid approach also improves interpretability, enabling models to correctly handle semantically misleading commits through more context-aware reasoning.